OpenStudy (anonymous):

compute the integral: (8x^2+6)/(x^2+1)(x+7)

OpenStudy (anonymous):

It is basically \[\int\limits_{}^{} 8x ^{2}+6 - \int\limits_{}^{}x ^{3}+7x ^{2}+x+7\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

i also have another question :( compute the indefinite integral: dx/(x^2+4)^(5/2)

OpenStudy (anonymous):

i am sorry, I was thinking of logarithms. Please ignore my previous answer. My integration is rusty. i'll get back to you with the correct answer.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the original question is; compute the integral:\[\int\limits_0{}^{1} (8x^2+6) / (x^2+1)(x+7) dx \]

OpenStudy (anonymous):

do you know integration by partial fractions? Understanding of that is essential to this problem http://www.math.ucdavis.edu/~kouba/CalcTwoDIRECTORY/partialfracdirectory/PartialFrac.html scroll down till you see how the partial fractions are done

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

This is how to solve your problem. The numerator in this example is also a quadratic equation and the denominator is a cubic equation: http://www.math.ucdavis.edu/~kouba/CalcTwoDIRECTORY/partialfracsoldirectory/PartialFracSol2.html#SOLUTION%209

OpenStudy (anonymous):

got it:)

OpenStudy (anonymous):

http://openstudy.com/updates/4d948f880ffe8b0bcdc9a720 can you help me with that problem ? :(